# Generating Exceptional Knots and Links with Arbitrary Braiding Topology

### 04 Mar 2026 v1

@@ Abstract
Non-Hermitian systems support band degeneracies fundamentally different from those in Hermitian systems, with exceptional points (EPs)—where eigenvalues and eigenvectors coalesce—standing out as defining singularities that have enabled transformative applications in optics, acoustics, circuits, and so on, from chiral mode switching and enhanced sensing to nonreciprocal transport. In three-dimensional (3D) non-Hermitian systems, such exceptional degeneracies extend into one-dimensional closed exceptional loops (ELs), whose global geometry can exhibit nontrivial knot and link topologies, a feature with no Hermitian analog and untapped potential for topological engineering. Here, we report a universal, constructive framework for realizing arbitrary knotted and linked ELs in 3D minimal two-band non-Hermitian systems, forging the first direct and systematic correspondence between knot theory and non-Hermitian band degeneracies. Leveraging Alexander’s theorem, we start from any knot or link encoded by a braid representation and design tight-binding Hamiltonians whose ELs faithfully replicate the prescribed knot/link topology in momentum space, enabling a rigorous classification of non-Hermitian topological phases via core knot invariants including braid words, Alexander polynomials, and Seifert surfaces. A key fundamental advance is that these knotted ELs are generically stable, requiring no symmetry protection or fine-tuning, and host non-Hermitian metallic phases characterized by Seifert surfaces that encode the defective nature of exceptional degeneracies—an essential distinction from Hermitian nodal lines, which rely on symmetries for their stability. We further demonstrate that knotted ELs can be continuously deformed and untied through deterministic topological transitions driven by a single tunable parameter, which induces controlled reconnection events of EPs and provides an experimentally actionable mechanism for manipulating global knot topology in momentum space. We also propose a concrete experimental realization of knotted and linked ELs in acoustic systems with tunable nonreciprocal couplings, validating the feasibility of observing these topological degeneracies in tabletop experiments. Parallel to the far-reaching impact of EPs across engineered platforms, our results establish knot and link topology as a universal organizing principle for non-Hermitian topological matter and unlock broad applicability for knotted ELs in photonic crystals, acoustic metamaterials, electrical circuits, and cold-atom systems—paving the way for next-generation topology-driven devices with reconfigurable topological modes, robust nonreciprocal responses, and novel spectral engineering capabilities.
@@
Exceptional degeneracies represent a defining hallmark of non-Hermitian physics, differing fundamentally from their Hermitian counterparts. At their core are exceptional points (EPs)—defective spectral singularities where eigenvalues and eigenvectors coalesce, rendering the Hamiltonian non-diagonalizable Heiss (2004); amd K. G. Makris et al. (2018); Ashida et al. (2020); Bergholtz et al. (2021). These degeneracies arise generically in open and driven systems and have been experimentally realized across diverse platforms, including photonics Feng et al. (2017); Pan et al. (2018); Miri and Alu (2019); Li et al. (2023); Feng et al. (2025); Xiao et al. (2025), acoustics Gao et al. (2021); Zhang et al. (2023a); Huang et al. (2024), mechanics Yoshida and Hatsugai (2019); Cui et al. (2023), electrical circuits Zhang and Franz (2020); Zhou et al. (2025), and active matter Sone et al. (2020, 2025). Acting as branch points on the complex energy Riemann surface, EPs give rise to striking physical effects such as chiral mode switching Li et al. (2020, 2025), mode selection Laha et al. (2020), nontrivial spectral winding Patil et al. (2022); Guria et al. (2024), and enhanced parametric response Parto et al. (2025); Zhou et al. (2025); Wang and Yan (2025), establishing them as fundamental building blocks of non-Hermitian phenomena.
@@
In momentum space, the dimensionality of exceptional degeneracies is governed by codimension arguments: in two-dimensional (2D) systems, EPs typically appear as isolated point defects Lin et al. (2019); Cao et al. (2024) or in pairs connected by open Fermi arcs Zhou et al. (2018a); Hu et al. (2025), while in three-dimensional (3D) systems, they generically extend into one-dimensional (1D) manifolds termed exceptional lines. These extended non-Hermitian degeneracies exhibit robustness against perturbations, enabling the formation of diverse configurations including closed exceptional rings (ERs) Xu et al. (2017); Yoshida and Hatsugai (2019), intersecting exceptional chains (ECs) Zhang et al. (2023b); Cui et al. (2023), and more elaborate geometries Zhang et al. (2024a). As embedded curves in 3D momentum space, exceptional lines possess an intrinsic global geometry that demands a topological characterization beyond mere local stability considerations.
@@
A compelling frontier in non-Hermitian band theory involves engineering nontrivial knot and link topologies within exceptional lines Carlström et al. (2019); Zhang et al. (2021b); Wang et al. (2021b); Carlström and Bergholtz (2018); Yang and Hu (2019); Zhang et al. (2020b); He et al. (2020). Knotted structures are fundamental to topology and pervade diverse physical contexts, from fluid vortices Machon and Alexander (2014), acoustic vortices Zhang et al. (2020a), and optical fields Dennis et al. (2010); Wang et al. (2025b, 2026) to polymers Zhang et al. (2025), liquid crystals Machon and Alexander (2014), quantum systems Hall et al. (2016), particle physics Eto et al. (2025), and topological solitons Veretenov et al. (2017). Embedding knotted structures into band degeneracies opens new avenues for non-Hermitian topological phases of matter, where abstract knot invariants may acquire direct physical significance through measurable spectral and dynamical signatures. Several prior works have demonstrated knotted or linked exceptional loops (ELs) in specific models Carlström et al. (2019); Zhang et al. (2021b); Wang et al. (2021b); Carlström and Bergholtz (2018); Yang and Hu (2019); Zhang et al. (2020b); He et al. (2020), revealing intriguing connections between non-Hermitian topology and exceptional structures.
@@
Beyond static topology, we further demonstrate that knotted ELs can be continuously untied through controlled topological transitions. Tuning a single continuous parameter that governs the effective width of the braid induces deterministic sequences of reconnection events, mediated by the motion, merging, and splitting of exceptional points (EPs). This process provides an experimentally accessible protocol for manipulating knot topology in momentum space, transforming knotted ELs into unlinked or unknotted configurations. Importantly, while the induced pathways are not necessarily minimal in the mathematical sense of the unknotting number, they are continuous, robust, and experimentally feasible—making them particularly well-suited for real-world implementations. More broadly, this work establishes a general bridge between knot theory and non-Hermitian band degeneracies, elevating ELs from geometric features to topological objects classified by knot invariants.
@@ II The defective Hamiltonian approach
The solutions to Eqs. (2) and the resulting geometry of the Fermi surface are strongly dependent on the dimensionality of the system. In 22D systems, each condition generally describes a 11D curve in the Brillouin zone. The intersection points of two curves are zero-dimensional (0D) degeneracies known as EPs. At these EPs, the Hamiltonian becomes defective, with coalescing eigenvalues and eigenvectors, leading to a branch point singularity in the complex energy Riemann surface. Generally, EPs in 2D systems appear in pairs and are connected by open Fermi arcs Zhou et al. (2018a); Hu et al. (2025).
@@ VII Summary and Discussions
The results presented in this work establish knotted and linked exceptional loops (ELs) as a broadly accessible and systematically designable class of non-Hermitian band degeneracies. By constructing a universal, direct mapping from braid representations to minimal two-band non-Hermitian Hamiltonians, our framework elevates knot and link topology from a model-dependent curiosity to a fundamental organizing principle for non-Hermitian band structures. This construction unifies all previously reported realizations of knotted and linked non-Hermitian degeneracies within a single theoretical pipeline Carlström et al. (2019); Zhang et al. (2021b); Wang et al. (2021b); Carlström and Bergholtz (2018); Yang and Hu (2019); Zhang et al. (2020b); He et al. (2020) and, crucially, demonstrates that arbitrary knot and link topologies can be engineered without symmetry protection or fine-tuning. This generic stability constitutes a defining distinction between non-Hermitian exceptional degeneracies and their Hermitian counterparts—nodal lines and nodal chains—whose existence and classification are inherently tied to symmetry constraints Burkov et al. (2011); Yan et al. (2017); Zhou et al. (2018b); Bzdušek et al. (2016); Yan et al. (2018).
@@
A central conceptual insight of this work is that the topology of ELs is encoded in the global geometry of the degeneracy manifold, rather than in local topological charges that characterize Hermitian band topology. In stark contrast to Hermitian nodal lines, whose stability and topological classification rely on rigid symmetry constraints Chiu et al. (2016); Po et al. (2017), ELs arise generically in 3D non-Hermitian systems as the intersection of real surfaces in momentum space. This intrinsic robustness frees EL topology from local symmetry requirements, enabling the realization of intricate knotted and linked configurations that are topologically forbidden or highly unnatural in Hermitian systems. From this perspective, classical knot invariants—including braid words, Alexander polynomials, and Seifert surfaces—acquire a direct physical interpretation as fundamental classifiers of non-Hermitian metallic phases, linking abstract topological mathematics to measurable spectral and structural properties of non-Hermitian matter. Our construction further unveils a deep, fundamental connection between non-Hermitian physics and classical knot theory, mediated by the use of semiholomorphic polynomials. These polynomials provide a natural, rigorous bridge between the algebraic description of knots and links in the three-sphere $\mathcal{S}^{3}$ and their physical realization as ELs in Bloch momentum space. While polynomial constructions for knotted curves have long been a subject of pure mathematical inquiry, their embedding into non-Hermitian Hamiltonians reveals a broader principle: non-Hermitian systems act as a fertile experimental and theoretical arena where abstract topological objects can be directly encoded into experimentally observable spectral features. In this sense, knotted and linked ELs are not merely geometric curiosities but tangible physical realizations of mathematical knots, endowed with distinct observable consequences—including spectral winding, defective eigenstates, and non-Hermitian open Fermi surfaces bounded by the EL degeneracy manifold.
@@
Beyond the static engineering of knotted EL topology, our demonstration of continuous untying of knotted ELs via a single tunable parameter introduces a new paradigm for topological manipulation in momentum space. Unlike Hermitian topological phase transitions, which are governed by gap closings at isolated high-symmetry points, the unknotting processes identified here proceed through deterministic, controlled reconnection events of exceptional points (EPs) and the transient generation of exceptional chains (ECs). ECs emerge as unavoidable intermediate states during EL untying, driving predictable morphological transformations of the degeneracy manifold without fine-tuning or symmetry breaking. These transitions establish a experimentally actionable route for altering global knot topology in momentum space, and they suggest that knot complexity itself can be treated as a tunable physical resource in non-Hermitian systems. An intriguing open question that emerges from this work is whether distinct unknotting pathways are characterized by universal spectral or dynamical features, or whether material-specific and experimental dynamical constraints select preferred topological transition routes in realistic non-Hermitian platforms.
@@
Several important and promising directions remain open for future exploration, building on the universal framework established here. First, while this work focuses on minimal two-band non-Hermitian models—the simplest platform for EL formation—extending the framework to multi-band non-Hermitian systems will likely unlock richer knotted and linked structures, enabled by higher-order exceptional degeneracies and nontrivial band permutations that have no two-band analogs. Second, the interplay between knotted ELs and additional symmetries—including crystalline, parity-time (PT), and non-Hermitian chiral symmetries—remains largely unexplored and may yield refined topological classification schemes, protected subclasses of knotted EL topology, or novel symmetry-enforced degeneracy transformations. Third, it would be of great fundamental and applied interest to investigate the dynamical and linear/nonlinear response phenomena associated with knotted ELs, including adiabatic encircling protocols Feilhauer et al. (2020); Arkhipov et al. (2023), nonreciprocal transport, non-Hermitian pumping Lee et al. (2020), and the signatures of knotted topology in wave-packet evolution and nonlinear spectral response Zheng et al. (2026).
@@
The universal framework developed herein extends far beyond the realm of non-Hermitian band theory, providing a general algorithm for engineering knotted topological structures that can serve as tailored initial conditions across diverse physical systems. The braid-based polynomial construction of knotted curves we introduce is broadly transferable, enabling the generation of knotted initial states for studying topological quantum dynamics in cold-atom and quantum optical systems Hall et al. (2016), the design of knotted optical fields with bespoke phase singularities for topological photonics Wang et al. (2025b, 2026), the investigation of knotted defect formation and evolution in liquid crystals and soft matter Ackerman and Smalyukh (2017); Hall et al. (2025), and the exploration of stable knotted solitons in classical and quantum field theories Amari and Sawado (2018). In this way, our work not only advances the understanding of non-Hermitian topological matter but also provides a cross-disciplinary tool for engineering knotted structures in physical science.
@@ Appendix B Knot invariants
B5 The relations between Seifert matrix and the Alexander polynomial

### 11 Mar 2026 v2

# Constructing Exceptional Knots and Links with Arbitrary Braiding Topology

@@ Abstract
Non-Hermitian systems support band degeneracies fundamentally different from those in Hermitian systems, with exceptional points (EPs)—where eigenvalues and eigenvectors coalesce—standing out as defining singularities that have enabled transformative applications in optics, acoustics, electrical circuits, and so on, from chiral mode switching and enhanced sensing to nonreciprocal transport. In three-dimensional (3D) non-Hermitian systems, such exceptional degeneracies extend into one-dimensional closed exceptional loops (ELs), whose global geometry can exhibit nontrivial knot and link topology, a feature with no Hermitian analog and untapped potential for topological engineering. Here, we report a universal construction framework for realizing exceptional knots and links with arbitrary braiding topology in 3D minimal two-band non-Hermitian systems, forging the first direct and systematic correspondence between knot theory and non-Hermitian band degeneracies. Leveraging Alexander’s theorem, we start from any knot or link encoded by a braid representation and design tight-binding Hamiltonians whose ELs faithfully replicate the prescribed knot/link topology in momentum space, enabling a rigorous classification of non-Hermitian topological phases via core knot invariants including braid words, Alexander polynomials, and Seifert surfaces. A key fundamental advance is that these knotted ELs are generically stable, requiring no symmetry protection or fine-tuning, and host non-Hermitian metallic phases characterized by Seifert surfaces that encode the defective nature of exceptional degeneracies—an essential distinction from Hermitian nodal lines, which rely on symmetries for their stability. We further demonstrate that knotted ELs can be continuously deformed and untied through deterministic topological transitions driven by a single tunable parameter, which induces controlled reconnection events of EPs and provides an experimentally actionable mechanism for manipulating global knot topology in momentum space. We also propose a concrete experimental realization of knotted and linked ELs in acoustic systems with tunable nonreciprocal couplings, validating the feasibility of observing these topological degeneracies in tabletop experiments. Parallel to the far-reaching impact of EPs across engineered platforms, our results establish knot and link topology as a universal organizing principle for non-Hermitian topological matter and unlock broad applicability for knotted ELs in photonic crystals, acoustic metamaterials, electrical circuits, and cold-atom systems—paving the way for next-generation topology-driven devices with reconfigurable topological modes, robust nonreciprocal responses, and novel spectral engineering capabilities.
@@
Exceptional degeneracies represent a defining hallmark of non-Hermitian physics, differing fundamentally from their Hermitian counterparts. At their core are exceptional points (EPs)—defective spectral singularities where eigenvalues and eigenvectors coalesce, rendering the Hamiltonian non-diagonalizable Heiss (2004); amd K. G. Makris et al. (2018); Ashida et al. (2020); Bergholtz et al. (2021). These degeneracies arise generically in open and driven systems and have been experimentally realized across diverse platforms, including photonics Feng et al. (2017); Pan et al. (2018); Miri and Alu (2019); Li et al. (2023); Feng et al. (2025); Xiao et al. (2025), acoustics Gao et al. (2021); Zhang et al. (2023a); Huang et al. (2024); Hu et al. (2025b), mechanics Yoshida and Hatsugai (2019); Cui et al. (2023), electrical circuits Zhang and Franz (2020); Zhou et al. (2025), and active matter Sone et al. (2020, 2025). Acting as branch points on the complex energy Riemann surface, EPs give rise to striking physical effects such as chiral mode switching Li et al. (2020, 2025), mode selection Laha et al. (2020), nontrivial spectral winding Patil et al. (2022); Guria et al. (2024), enhanced parametric response Parto et al. (2025); Zhou et al. (2025); Wang and Yan (2025) and unidirectional amplification Xue et al. (2021); Wen et al. (2022), establishing them as fundamental building blocks of non-Hermitian phenomena.
@@
In momentum space, the dimensionality of exceptional degeneracies is governed by codimension arguments: in two-dimensional (2D) systems, EPs typically appear as isolated point defects Lin et al. (2019); Hu et al. (2022); Cao et al. (2024) or in pairs connected by open Fermi arcs Zhou et al. (2018a); Yang et al. (2021); Hu et al. (2025a), while in three-dimensional (3D) systems, they generically extend into one-dimensional (1D) manifolds termed exceptional lines. These extended non-Hermitian degeneracies exhibit robustness against perturbations, enabling the formation of diverse configurations including closed exceptional rings (ERs) Xu et al. (2017); Yoshida and Hatsugai (2019); Hu et al. (2025c) and high-order ERs Hu et al. (2025b), intersecting exceptional chains (ECs) Zhang et al. (2023b); Cui et al. (2023), and more elaborate geometries Zhang et al. (2024a). As embedded curves in 3D momentum space, exceptional lines possess an intrinsic global geometry that demands a topological characterization beyond mere local stability considerations.
@@
A compelling frontier in non-Hermitian band theory involves engineering nontrivial knot and link topologies within exceptional lines Carlström et al. (2019); Zhang et al. (2021b); Wang et al. (2021b); Carlström and Bergholtz (2018); Yang and Hu (2019); Zhang et al. (2020b); He et al. (2020). Knotted structures are fundamental to topology and pervade diverse physical contexts, from fluid vortices Machon and Alexander (2014), acoustic vortices Zhang et al. (2020a), and optical fields Dennis et al. (2010); Wang et al. (2025c, 2026) to polymers Zhang et al. (2025), liquid crystals Machon and Alexander (2014), quantum systems Hall et al. (2016), particle physics Eto et al. (2025), and topological solitons Veretenov et al. (2017). Embedding knotted structures into band degeneracies opens new avenues for non-Hermitian topological phases of matter, where abstract knot invariants may acquire direct physical significance through measurable spectral and dynamical signatures. Several prior works have demonstrated knotted or linked exceptional loops (ELs) in specific models Carlström et al. (2019); Zhang et al. (2021b); Wang et al. (2021b); Carlström and Bergholtz (2018); Yang and Hu (2019); Zhang et al. (2020b); He et al. (2020), revealing intriguing connections between non-Hermitian topology and exceptional structures.
@@
Beyond static topology, we further demonstrate that knotted ELs can be continuously untied through controlled topological transitions. Tuning a single continuous parameter that governs the effective width of the braid induces deterministic sequences of reconnection events, mediated by the motion, merging, and splitting of EPs. This process provides an experimentally accessible protocol for manipulating knot topology in momentum space, transforming knotted ELs into unlinked or unknotted configurations. Importantly, while the induced pathways are not necessarily minimal in the mathematical sense of the unknotting number, they are continuous, robust, and experimentally feasible—making them particularly well-suited for real-world implementations. More broadly, this work establishes a general bridge between knot theory and non-Hermitian band degeneracies, elevating ELs from geometric features to topological objects classified by knot invariants.
@@ II The defective Hamiltonian approach
The solutions to Eqs. (2) and the resulting geometry of the Fermi surface are strongly dependent on the dimensionality of the system. In 22D systems, each condition generally describes a 11D curve in the Brillouin zone. The intersection points of two curves are zero-dimensional (0D) degeneracies known as EPs. At these EPs, the Hamiltonian becomes defective, with coalescing eigenvalues and eigenvectors, leading to a branch point singularity in the complex energy Riemann surface. Generally, EPs in 2D systems appear in pairs and are connected by open Fermi arcs Zhou et al. (2018a); Hu et al. (2025a).
@@ VII Summary and Discussions
The results presented in this work establish knotted and linked ELs as a broadly accessible and systematically designable class of non-Hermitian band degeneracies. By constructing a universal, direct mapping from braid representations to minimal two-band non-Hermitian Hamiltonians, our framework elevates knot and link topology from a model-dependent curiosity to a fundamental organizing principle for non-Hermitian band structures. This construction unifies all previously reported realizations of knotted and linked non-Hermitian degeneracies within a single theoretical pipeline Carlström et al. (2019); Zhang et al. (2021b); Wang et al. (2021b); Carlström and Bergholtz (2018); Yang and Hu (2019); Zhang et al. (2020b); He et al. (2020) and, crucially, demonstrates that arbitrary knot and link topologies can be engineered without symmetry protection or fine-tuning. This generic stability constitutes a defining distinction between non-Hermitian exceptional degeneracies and their Hermitian counterparts—nodal lines and nodal chains—whose existence and classification are inherently tied to symmetry constraints Burkov et al. (2011); Yan et al. (2017); Zhou et al. (2018b); Bzdušek et al. (2016); Yan et al. (2018).
@@
A central conceptual insight of this work is that the topology of ELs is encoded in the global geometry of the degeneracy manifold, rather than in local topological charges that characterize Hermitian band topology. In stark contrast to Hermitian nodal lines, whose stability and topological classification rely on rigid symmetry constraints Chiu et al. (2016); Po et al. (2017); Kruthoff et al. (2017), ELs arise generically in 3D non-Hermitian systems as the intersection of real surfaces in momentum space. This intrinsic robustness frees EL topology from local symmetry requirements, enabling the realization of intricate knotted and linked configurations that are topologically forbidden or highly unnatural in Hermitian systems. From this perspective, classical knot invariants—including braid words, Alexander polynomials, and Seifert surfaces—acquire a direct physical interpretation as fundamental classifiers of non-Hermitian metallic phases, linking abstract topological mathematics to measurable spectral and structural properties of non-Hermitian matter. Our construction further unveils a deep, fundamental connection between non-Hermitian physics and classical knot theory, mediated by the use of semiholomorphic polynomials. These polynomials provide a natural, rigorous bridge between the algebraic description of knots and links in the three-sphere $\mathcal{S}^{3}$ and their physical realization as ELs in Bloch momentum space. While polynomial constructions for knotted curves have long been a subject of pure mathematical inquiry, their embedding into non-Hermitian Hamiltonians reveals a broader principle: non-Hermitian systems act as a fertile experimental and theoretical arena where abstract topological objects can be directly encoded into experimentally observable spectral features. In this sense, knotted and linked ELs are not merely geometric curiosities but tangible physical realizations of mathematical knots, endowed with distinct observable consequences—including spectral winding, defective eigenstates, and non-Hermitian open Fermi surfaces bounded by the EL degeneracy manifold.
@@
Beyond the static engineering of knotted EL topology, our demonstration of continuous untying of knotted ELs via a single tunable parameter introduces a new paradigm for topological manipulation in momentum space. Unlike Hermitian topological phase transitions, which are governed by gap closings at isolated high-symmetry points, the unknotting processes identified here proceed through deterministic, controlled reconnection events of EPs and the transient generation of ECs. ECs emerge as unavoidable intermediate states during EL untying, driving predictable morphological transformations of the degeneracy manifold without fine-tuning or symmetry breaking. These transitions establish a experimentally actionable route for altering global knot topology in momentum space, and they suggest that knot complexity itself can be treated as a tunable physical resource in non-Hermitian systems. An intriguing open question that emerges from this work is whether distinct unknotting pathways are characterized by universal spectral or dynamical features, or whether material-specific and experimental dynamical constraints select preferred topological transition routes in realistic non-Hermitian platforms.
@@
Several important and promising directions remain open for future exploration, building on the universal framework established here. First, while this work focuses on minimal two-band non-Hermitian models—the simplest platform for EL formation—extending the framework to multi-band non-Hermitian systems will likely unlock richer knotted and linked structures, enabled by higher-order exceptional degeneracies and nontrivial band permutations that have no two-band analogs. Second, the interplay between knotted ELs and additional symmetries—including crystalline, parity-time (PT), and non-Hermitian chiral symmetries—remains largely unexplored and may yield refined topological classification schemes, protected subclasses of knotted EL topology, or novel symmetry-enforced degeneracy transformations. Third, it would be of great fundamental and applied interest to investigate the dynamical and linear/nonlinear response phenomena associated with knotted ELs, including adiabatic encircling protocols Feilhauer et al. (2020); Arkhipov et al. (2023), nonreciprocal transport, non-Hermitian pumping Lee et al. (2020), and the signatures of knotted topology in wave-packet evolution Wang et al. (2025b) and nonlinear spectral response Zheng et al. (2026).
@@
The universal framework developed herein extends far beyond the realm of non-Hermitian band theory, providing a general algorithm for engineering knotted topological structures that can serve as tailored initial conditions across diverse physical systems. The braid-based polynomial construction of knotted curves we introduce is broadly transferable, enabling the generation of knotted initial states for studying topological quantum dynamics in cold-atom and quantum optical systems Hall et al. (2016), the design of knotted optical fields with bespoke phase singularities for topological photonics Wang et al. (2025c, 2026), the investigation of knotted defect formation and evolution in liquid crystals and soft matter Ackerman and Smalyukh (2017); Hall et al. (2025), and the exploration of stable knotted solitons in classical and quantum field theories Amari and Sawado (2018). In this way, our work not only advances the understanding of non-Hermitian topological matter but also provides a cross-disciplinary tool for engineering knotted structures in physical science.
@@ Appendix B Knot invariants
B5 Relations between Seifert matrix and the Alexander polynomial